Find the uncertainity in the position of an electron (mass 9.1 x 10^-28) moving with a velocity of 3 x 10^4 cm sec^-1, accurate upto 0.011%
 

Grade:11

1 Answers

Vikas TU
4 years ago
Hiiii 
According t Heisenberg uncertainity principle te product of uncertainity in position delta x and delta v > h/n*pie
Uncertainity in velocity = 300*0.011/100 = 0.033 m/sec
Delta x = h/(4*pie*m*delta u )
put the values you will gwt delta x as 0.1755 m